OpenStudy (anonymous):

write the equation -2x=-4y+12 in slope-intecept form label the slope and y intercept.

OpenStudy (anonymous):

for slope i get 1/2 and for y intecept i get 15 is this correct?

OpenStudy (anonymous):

-2x=-4y+12 so, 4y=2x+12 y=1/2x+3 so be happy :) you are right

OpenStudy (anonymous):

Thank you